OpenStudy (missmeow):

Am I doing this right? Which of these is a simplified form of the equation 7x + 4 = 5 + 3x + 1x? A. 3x = 9 B. 3x = 1 C. 11x = 9 D. 11x = 1 My process: 7x + 4 = 5 + 3x + 1x 7x + 4 = 5 + 4x 7x + 4x = 5 + 4 11x + 9 So, I want to say C. Am I right?

OpenStudy (freckles):

let's look at your 7x+4=5+4x you are right here but there are incorrect steps take afterwards

OpenStudy (missmeow):

What did I do wrong?

OpenStudy (freckles):

you know that 4x-4x=0 and that 4-4=0 right?

OpenStudy (anonymous):

you are on the right track. Except when you put terms on another side, you must do the opposite

OpenStudy (freckles):

You have used that 4x+4x=0 and that 4+4=0 when this is not true

OpenStudy (freckles):

7x+4=5+4x Pretend we want to get the x's on all one side right now what do you think we must do to do this

OpenStudy (missmeow):

I used the distribution property to move the 4 in the 4x's place.

OpenStudy (anonymous):